# 1D-CNN Optimization for Non-contact Respiration Pattern Classification

### v1

@@ Abstract
Index Terms:
Respiration classification, Light-wave sensing, Infrared sensing, Non-contact sensing, Genetic algorithm, Transfer learning
I Introduction
Human respiratory rate and its pattern convey valuable information about the physical and mental states of the subject. Anomalous breathing can be indicative of serious health issues such as asthma, obstructive sleep apnea, bronchitis, emphysema, lung cancer, chronic obstructive pulmonary disease, COVID-19, etc. It can also serve as a precursor to unstable mental conditions, including stress, panic, anxiety, fatigue, anger, etc. Contact-based monitoring falls short in capturing the true essence of one’s breathing, as individuals may alter their breathing patterns when aware of being monitored. Furthermore, subjects may be too young or too ill (such as patients with contagious diseases or those in burn units) to apply contact-based approaches for an extended duration. Current automated technologies for non-contact breathing monitoring typically rely on sensing electromagnetic signals, such as radar [1, 2, 3, 4], WiFi [5, 6, 7, 8], or extracting breathing information from videos captured by red-green-blue (RGB) [9, 10, 11, 12, 13] or thermal infrared cameras [14, 15, 16, 17]. We have developed a non-contact respiration monitoring system based on sensing incoherent infrared (IR) light reflected from the subject’s chest. This technology surpasses existing methods due to the safe, ubiquitous, and discreet nature of infrared light, avoiding privacy issues associated with camera-based approaches. The main functional components of the system include the LWS hardware, signal processing, and feature extraction algorithms, along with classification algorithms. Infrared LED sources illuminate the subject’s chest, and the reflected light’s intensity, varying with chest movement, is captured and converted into electrical voltage by a commercially available photodetector. Collected data are processed to extract features using signal processing algorithms and then classified using machine learning models.
@@ I Introduction
The remainder of this manuscript is organized as follows. Section II describes the system model including hardware setup and software algorithms to collect breathing data of different classes and perform classification. Section III presents the details of 1-dimensional convolutional neural network (1D-CNN) optimization including the application of transfer learning in subsection III-A and genetic algorithm (GA) in subsection III-B. Next, Section IV presents the results and discussion from GA and final breathing classification using the obtained optimized model. Finally, Section V draws conclusions from the effort and forecasts future research directions.
II System Design and Implementation
II-A Experimental setup
Human subjects typically cannot replicate breathing with precise frequency, amplitude, and patterns consistently. Therefore, a mechanical robot capable of simulating human breathing in various programmable patterns was developed and utilized for this study. The LWS sensing hardware includes infrared light sources as transmitters, a photodetector as the receiver, a digital signal processing (DSP) unit to convert the received signal from analog to digital, and an electronic module for processing and storing the digital data. Additionally, a function generator and lock-in amplifier were used to for coherent detection of the received light. The light source utilized was an invisible 940 nm IR lamp board (48 black LED illuminator array), offering a range of 30 ft and a wide beamwidth of 120° wide beamwidth [27]. The light source was modulated by a 1 kHz sinusoidal voltage wave from a function generator. The receiver employed was a commercial photodetector, Thorlabs PDA100A [28] with a converging lens. For lock-in detection, SR830DSP frequency lock-in amplifier [29] was utilized. A Raspberry Pi, equipped with a PiPlate analog-to-digital converter (ADC) circuit [30], managed the data collection, digitization, and storage for subsequent offline processing. The complete experimental setup is illustrated in Fig. 1.
TABLE I: Characteristics and number of collected data of each breathing class
@@ II-A Experimental setup
0
Eupnea
12-20
30-58
300
1
Apnea
0
0
300
2
Tachypnea
21-50
30-58
300
3
Bradypnea
1-11
30-58
300
4
Hyperpnea
12-20
59-100
300
5
Hypopnea
12-20
1-29
300
6
Kussmaul’s
21-50
59-100
300
7
Faulty data
Any
Any
300
Total
2400
II-B Data collection
To gather respiration data, the infrared light emitted by the LED source underwent modulation using a sinusoidal voltage wave at 1 kHz with a peak-to-peak amplitude of 3.8 V and a DC offset of 8.1 V from the function generator. These specific voltage parameters were chosen to ensure linear LED operation with sufficiently high light intensity. The reflected light was captured by a photodetector paired with a converging lens having a focal length of 25.4 mm. The photodetector’s gain was set to 40 dB. The output from the photodetector was linked to a lock-in amplifier with a time constant of 100 ms. The resulting filtered signal from the lock-in amplifier was then digitized by the ADC. Finally, a Python script executed on the Raspberry Pi collected and stored the voltage amplitude data, along with their corresponding timestamps.
@@ II-B Data collection
From the literature, 7 different classes of human breathing were identified and their characteristics, in terms of breathing rate (frequency) and depth (amplitude) [^1] Breathing depth is measured from the human rib cage movement and it is expressed as a percentage of the maximum movement of the rib cage [31]., were summarized in Table I [32, 4, 31, 33, 34, 14, 35, 36, 37, 38, 39, 40]. One of them (Eupnea) was normal breathing, while the rest were examples of different categories of abnormal breathing. The robot was programmed to breathe with frequencies and amplitudes in the appropriate ranges according to Table I to generate data experimentally for each class. Data for one additional class called ‘faulty data’ were generated too by manually interrupting the system or the surroundings during data collection. This class was introduced to detect erroneous data generated due to possible internal or external malfunction so that they can be discarded and recollected. $\sin^{6}$ waveform seemed to have the closest match with human breathing in terms of inspiration and expiration time, hence this pattern was used in the robot for all breathing classes. A total of 2400 data instances (300 per class) were collected during the daytime, keeping the windows of the room unshaded and internal lighting common for an office environment. For each class, 100 data instances were collected at 0.5 m, 1 m and 1.5 m distances each. Each data was 30 s long and collected at 100 Hz sampling frequency. All the collected data were saved along with their class labels (0 to 7) for further offline processing. Fig. 2 shows the time domain waveforms of a few samples of raw data, one from each class, collected at 1 m distance.
@@
Figure 5: Genetic algorithm workflow.
III 1D-CNN Optimization
III-A Transfer Learning Application
Transfer learning is a powerful paradigm in machine learning that leverages knowledge gained from solving one task to improve the performance on a different, but related, task. In essence, it allows a model trained on one dataset to be fine-tuned or adapted for a different but similar task. This approach is particularly beneficial when the target task has limited labeled data, as it enables the model to generalize better. It is also useful to make the training faster through utilizing pre-trained weights. In the context of deep learning, transfer learning is often applied using pre-trained neural network architectures, such as those trained on large-scale image datasets like ImageNet [41]. The model’s initial layers learn generic features like edges and textures, while the deeper layers capture more task-specific information. This modular structure facilitates the reuse of the pre-trained layers for a new task. For instance, in Python, utilizing transfer learning with TensorFlow library involves loading pre-trained models like VGG16 [42] or ResNet [43] and adapting them for specific tasks with minimal effort. This accelerates training and enhances performance, making transfer learning a valuable tool in the machine learning toolbox.
Although many famous 2D-CNN models pre-trained on ImageNet dataset are available as API for direct use, their 1D counterparts are rare in the literature and GitHub repositories [20]. To address this problem, a single 1D-CNN model was obtained through random search that yields higher classification accuracy than most of the other tested models. The complete architecture of that model is shown in the leftmost bounded box in Fig. 3 and described as follows:
@@ III-A Transfer Learning Application
III-B Optimization Using Genetic Algorithm
Genetic algorithm was applied using Python to find the optimum 1D-CNN model from a chosen solution space through crossovers and mutations over generations of solutions. Since transfer learning was applied to preload part of the CNN model in trained status, GA was used mainly to tune the rest of the model. The high level workflow of the genetic algorithm applied was shown in Fig. 5, but it involved some advanced features too which will be discussed here.
@@ III-B Optimization Using Genetic Algorithm
III-B1 Build Chromosome and Initial Population
The first step was to decide the chromosome structure or genotypes of each gene and define their corresponding phenotypes or how they relate to the 1D-CNN architecture. Our chromosome contains four genes as shown in Fig. 6. We used all parameters as powers of 2 because these numbers create more computationally efficient model. First gene denotesd the number of kernels for the newly added convolutional layer and is a random integer number between 3 to 8. If the integer generated is 6, then the number of kernels for that layer will be $642^{6}=$ 64. Second gene denotes the length of kernels and is an integer between 2 to 6, or equivalently kernel length varied between $2^{3}$ and $2^{6}$. The third gene represents the maxpooling size and the fourth gene indicates the number of neurons present in the fully connected layer.
III-B2 Get Fitness Values
@@
Figure 6: Chromosome for GA.
III-B3 Perform Crossover
@@ III-B5 Perform Selection
IV Results and Discussions
The results from a few trial runs of GA up to 200 generations are displayed in Fig. 7. It shows how the maximum fitness and the average fitness evolves over generations. 2 trials are shown with highest fitness individuals selected as crossover parents, while 2 more are shown with Roulette wheel selection for selecting parents for crossover. Average fitness values fluctuate frequently, hence 10 points moving average filtering was performed to see the trend. Although the average fitness values are slowly increasing, there is clear increasing trend in the maximum fitness values. Trial 1 with the Roulette wheel parents selection for crossover yielded the highest fitness value 0.43 or 43% classification accuracy after 1 epoch. The corresponding chromosome was [7, 5, 1, 8] and this was chosen as the best solution to test on the entire dataset over many epochs. This solution was combined with the pre-trained part of the model and then the entire model was trained on the available 2400 data instances (without augmentation this time) for 30 epochs. Two distinct training approaches were employed: one involving training the model from scratch with no pre-trained weights, and the other incorporating transfer learning. The latter utilized pre-trained weights for the initial layers, while training only the remaining layers. The training and test accuracies as a function of number of epochs for both scenarios are plotted in Fig. 8. TNotably, the achieved test accuracies were comparable in both training methods, with a slightly higher accuracy (87.33%, compared to 86.33%) observed when incorporating transfer learning in the final training.
V Conclusion and Future Directions

### v2

@@ Abstract
Keywords:
Respiration monitoring Light-wave sensing Evolutionary optimization Genetic algorithm Transfer learning.
1 Introduction
Human respiratory rate and its pattern convey valuable information about the physical and mental states of the subject. Anomalous breathing can be indicative of serious health issues such as asthma, bronchitis, emphysema, lung cancer, COVID-19, etc. It can also serve as a precursor to unstable mental conditions, including stress, panic, anxiety, etc. Contact-based monitoring falls short in capturing the true essence of one’s breathing, as individuals may alter their breathing patterns when aware of being monitored. Furthermore, subjects may be too young or too ill to apply contact-based approaches for an extended duration. Current automated technologies for non-contact breathing monitoring typically rely on sensing electromagnetic signals, such as radar, WiFi, or extracting breathing information from videos captured by red-green-blue (RGB) or thermal infrared cameras [1, 2, 3, 4, 5]. We have developed a non-contact respiration monitoring system based on sensing incoherent infrared (IR) light reflected from the subject’s chest. This technology surpasses existing methods due to the safe, ubiquitous, and discreet nature of infrared light, avoiding privacy issues associated with camera-based approaches. The main functional components of the system include the LWS hardware, signal processing, and feature extraction algorithms, along with classification algorithms. Infrared light-emitting diode (LED) sources illuminate the subject’s chest, and the reflected light’s intensity, varying with chest movement, is captured and converted into electrical voltage by a commercially available photodetector. Collected data are processed to extract features using signal processing algorithms and then classified using machine learning models.
@@ 1 Introduction
The remainder of this manuscript is organized as follows. Section 2 describes the system model including hardware setup and software algorithms to collect breathing data of different classes and perform classification. Section 3 presents the details of 1D-CNN optimization including the application of transfer learning and genetic algorithm. Next, Section 4 presents the results and discussion from GA and final breathing classification using the obtained optimized model. Finally, Section 5 draws conclusions from the effort and forecasts future research directions.
2 System Design and Implementation
2.1 Experimental setup
Human subjects typically cannot replicate breathing with precise frequency, amplitude, and patterns consistently. Therefore, a mechanical robot capable of simulating human breathing in various programmable patterns was developed and utilized for this study. The LWS sensing hardware includes infrared light sources as transmitters, a photodetector as the receiver, a digital signal processing (DSP) unit to convert the received signal from analog to digital, and an electronic module for processing and storing the digital data. Additionally, a function generator and lock-in amplifier were used to for coherent detection of the received light. The light source utilized was an invisible 940 nm IR lamp board (48 black LED illuminator array), offering a range of 30 ft and a wide beamwidth of 120° wide beamwidth. The light source was modulated by a 1 kHz sinusoidal voltage wave from a function generator. The receiver employed was a commercial photodetector, Thorlabs PDA100A with a converging lens. For lock-in detection, SR830DSP frequency lock-in amplifier was utilized. A Raspberry Pi, equipped with an analog-to-digital converter (ADC) circuit, managed the data collection, digitization, and storage for subsequent offline processing. The complete experimental setup is illustrated in Fig. 1.
Table 1: Characteristics and number of collected data of each breathing class
@@ 2.1 Experimental setup
0 Eupnea 12-20 30-58 300
1 Apnea 0 0 300
2 Tachypnea 21-50 30-58 300
3 Bradypnea 1-11 30-58 300
4 Hyperpnea 12-20 59-100 300
5 Hypopnea 12-20 1-29 300
6 Kussmaul’s 21-50 59-100 300
7 Faulty data Any Any 300
Total 2400
2.2 Data collection
To gather respiration data, the infrared light emitted by the LED source underwent modulation using a sinusoidal voltage wave of 1 kHz frequency from the function generator. The reflected light was captured by a photodetector paired with a converging lens. The photodetector’s gain was set to 40 dB. The output from the photodetector was linked to a lock-in amplifier with a time constant of 100 ms. The resulting filtered signal from the lock-in amplifier was then digitized by the ADC. Finally, a Python script executed on the Raspberry Pi collected and stored the voltage amplitude data, along with their corresponding timestamps.
@@ 2.2 Data collection
From the literature, 7 different classes of human breathing were identified and their characteristics, in terms of breathing rate (frequency) and depth (amplitude) [^1] Breathing depth is measured from the human rib cage movement and it is expressed as a percentage of the maximum movement of the rib cage [15]., were summarized in Table 1 [16, 1, 15, 17, 18, 4, 19, 20, 21]. One of them (Eupnea) was normal breathing, while the rest were examples of different categories of abnormal breathing. The robot was programmed to breathe with frequencies and amplitudes in the appropriate ranges according to Table I to generate data experimentally for each class. Data for one additional class called ‘faulty data’ were generated too by manually interrupting the system or the surroundings during data collection. This class was introduced to detect erroneous data generated due to possible internal or external malfunction so that they can be discarded and recollected. $\sin^{6}$ waveform seemed to have the closest match with human breathing in terms of inspiration and expiration time, hence this pattern was used in the robot for all breathing classes. A total of 2400 data instances (300 per class) were collected during the daytime, keeping the windows of the room unshaded and internal lighting common for an office environment. For each class, 100 data instances were collected at 0.5 m, 1 m and 1.5 m distances each. Each data was 30 s long and collected at 100 Hz sampling frequency. All the collected data were saved along with their class labels (0 to 7) for further offline processing. Fig. 2 shows the time domain waveforms of a few samples of raw data, one from each class, collected at 1 m distance.
@@
3 1D-CNN Optimization
3.1 Transfer Learning Application
Transfer learning is a powerful paradigm in machine learning that leverages knowledge gained from solving one task to improve the performance on a different, but related, task. In essence, it allows a model trained on one dataset to be fine-tuned or adapted for a different but similar task. This approach is particularly beneficial when the target task has limited labeled data, as it enables the model to generalize better. It is also useful to make the training faster through utilizing pre-trained weights. In the context of deep learning, transfer learning is often applied using pre-trained neural network architectures, such as those trained on large-scale image datasets like ImageNet. The model’s initial layers learn generic features like edges and textures, while the deeper layers capture more task-specific information. This modular structure facilitates the reuse of the pre-trained layers for a new task. For instance, in Python, utilizing transfer learning with TensorFlow library involves loading pre-trained models like VGG16 or ResNet and adapting them for specific tasks with minimal effort. This accelerates training and enhances performance, making transfer learning a valuable tool in the machine learning toolbox.
@@ 3.1 Transfer Learning Application
Although many famous 2D-CNN models pre-trained on ImageNet dataset are available as API for direct use, their 1D counterparts are rare in the literature and GitHub repositories. To address this problem, a single 1D-CNN model was obtained through random search that yields higher classification accuracy than most of the other tested models. The complete architecture of that model is shown in the leftmost bounded box in Fig. 3 and described as follows:
@@
Figure 6: Chromosome for GA.
3.2 Optimization Using Genetic Algorithm
Genetic algorithm was applied using Python to find the optimum 1D-CNN model from a chosen solution space through crossovers and mutations over generations of solutions. Since transfer learning was applied to preload part of the CNN model in trained status, GA was used mainly to tune the rest of the model. The high-level workflow of the genetic algorithm applied was shown in Fig. 5, but it involved some advanced features too which will be discussed here.
@@ 3.2 Optimization Using Genetic Algorithm
3.2.1 Build Chromosome and Initial Population:
The first step was to decide the chromosome structure or genotypes of each gene and define their corresponding phenotypes or how they relate to the 1D-CNN architecture. Our chromosome contains four genes as shown in Fig. 6. We used all parameters as powers of 2 because these numbers create more computationally efficient model. First gene denotes the number of kernels for the newly added convolutional layer and is a random integer number between 3 to 8. If the integer generated is 6, then the number of kernels for that layer will be $642^{6}=$ 64. Second gene denotes the length of kernels and is an integer between 2 to 6, or equivalently kernel length varied between $2^{3}$ and $2^{6}$. The third gene represents the maxpooling size and the fourth gene indicates the number of neurons present in the fully connected layer.
3.2.2 Get Fitness Values:
@@
3.2.3 Perform Crossover:
@@ 3.2.5 Perform Selection:
4 Results and Discussion
The results from a few trial runs of GA up to 200 generations are displayed in Fig. 7. It shows how the maximum fitness and the average fitness evolves over generations. 2 trials are shown with highest fitness individuals selected as crossover parents, while 2 more are shown with Roulette wheel selection for selecting parents for crossover. Average fitness values fluctuate frequently, hence 10 points moving average filtering was performed to see the trend. Although the average fitness values are slowly increasing, there is clear increasing trend in the maximum fitness values. Trial 1 with the Roulette wheel parents selection for crossover yielded the highest fitness value 0.43 or 43% classification accuracy after 1 epoch. The corresponding chromosome was [7, 5, 1, 8] and this was chosen as the best solution to test on the entire dataset over many epochs. This solution was combined with the pre-trained part of the model and then the entire model was trained on the available 2400 data instances (without augmentation this time) for 30 epochs. Two distinct training approaches were employed: one involving training the model from scratch with no pre-trained weights, and the other incorporating transfer learning. The latter utilized pre-trained weights for the initial layers, while training only the remaining layers. The training and test accuracies as a function of number of epochs for both scenarios are plotted in Fig. 8. Notably, the achieved test accuracies were comparable in both training methods, with a slightly higher accuracy (87.33%, compared to 86.33%) observed when incorporating transfer learning in the final training.
5 Conclusion and Future Directions
@@
5.0.1 Acknowledgements
This work was supported in part by the U.S. National Science Foundation under Grants 2008556, 2323301 and 2336852.